# Hyperuniformity and optimal transport of point processes

### v3

@@ 1 Introduction
The celebrated AKT (Ajtai, Komlós and Tusnády) theorem [1] states that, if one draws $2n$ i.i.d. uniform points $X_{1},\dots,X_{n},Y_{1},\dots,Y_{n}$ in a two-dimensional cube with volume $n$, it is possible to pair each $X_{i}$ with an exclusive $Y_{\sigma(i)}$ for some permutation $\sigma$ such that the typical length $|X_{1}-Y_{\sigma(1)}|$ is of the order $\sqrt{\ln(n)}$ with high probability, and it is the best possible rate. The role of the dimension $2$ is very important, as for the same statement in dimension $1$, the typical distance is of order $\sqrt{n}$, and in dimension $d\geqslant 3$ it is of order $1$ This result has known many refinements and generalisations such as replacing the total distance by weighted versions $\sum|X_{i}-Y_{\sigma(i)}|^{p},p>0$, drawing points with a non-uniform law on the cube, giving large deviations [23], computing exact asymptotic constants [4, 31] etc. Using the framework of optimal transport, these asymptotics are equivalent to the cost of optimal transport between the measure $\nu=\sum_{i=1}^{n}\delta_{Y_{i}}$ and the measure $\mu=\sum_{i=1}^{n}\delta_{X_{i}}$ on the cube. We refer the reader to [52, 56, 43] for surveys of such results and for a sampler of other recent developments, see [6, 13, 26].
@@
$\mathbf{E}\left(\frac{X^{d/2}}{1+|\ln(X)|^{\gamma}}\right)<\infty\textrm{ for $d=1,2$, any $\gamma>1$ and }\mathbf{E}(X^{2})<\infty\textrm{ for $d\geqslant 3$},$ (1)
@@
The starting point of this article is to replace $\mu$ and/or $\nu$ by arbitrary stationary random point processes and consider matchings between $\mu$ and $\nu$ and equivalently transport map or plan between the Lebesgue measure and $\mu$. In specific planar cases such as zeros of Gaussian entire functions (GEF) or Ginibre point process, existence of good matchings have been shown. In particular, [53] shows exponential moments for the matching cost between zeros of GEF and lattice and [35, 49] prove respectively finite first and second moments for the transport of Ginibre ensemble to the Lebesgue measure. Both these point processes are known to be hyperuniform (also known as super-homogeneous) i.e., the asymptotic number variance (i.e. variance of the number of points) on a large window grows more slowly than volume order, unlike the Poisson point process [58, 24]. Hyperuniform (HU) processes appear in various contexts such as statistical physics, biology, material sciences, random matrices, dynamical systems, numerical integration [57, 20]. A concise description coined by physicists about HU processes is “global order and local disorder”, meaning that even though the particles are locally disordered, exhibit isotropy and asymptotic independence, they should behave at large scales as lattices, and should in some sense be arranged in a nice orderly manner. A natural question is therefore whether such processes are well distributed in space, and we believe that this should be reflected by their transport properties, in the sense that a nicely spread point sample should be easy to transport to the Lebesgue measure, or equivalently to a regular lattice. Notable examples of hyperuniform point processes are constructed as a good matching to a lattice; see [48, 39, 44].
@@
We see from (1) that this property (existence of a matching with finite second moment for typical cost) is true in dimension $d\geqslant 3$ for Poisson processes, and we will see that it extends to other processes, whether HU or not, as long as they are not hyperfluctuating. In dimension $2$, we answer the question on existence of ’good’ matchings for HU processes in the affirmative. In other words, we show that for a HU process $\mu$ in dimension $d=2$, identifying $\mu$ with its support, under mild logarithmic integrability conditions on the reduced pair correlation function, there is a $L^{2}$ process $\{T(k);k\in\mathbb{Z}^{2}\}$ invariant under $\mathbb{Z}^{2}$ translations such that $\mu=\sum_{k\in\mathbb{Z}^{2}}\delta_{k+T(k)}$, meaning it is a $L^{2}$ perturbed lattice. Many known HU processes satisfy this condition, including the Ginibre point process and zeros of Gaussian entire function. An example of a HU process with infinite $L^{1}$-matching has been provided in [32, Theorem 2(2)] indicating the necessity of some additional integrability condition but whether logarithmic integrability is required or not is moot. Our interest in $L^{2}$ perturbed lattices is also in part due to the following recent results on perturbed lattices due to [21]. Namely it was shown that $L^{2}$ perturbed lattices are HU but for any $\delta>0$, there exist $L^{2-\delta}$ perturbed lattices which are not hyperuniform.
@@
En route to our result for HU processes, we generalize the above bounds (1) for Poisson processes to stationary random point processes with integrable reduced pair correlation function (RPCM). This is to say that in magnitude, the matching bounds for point processes with number variance growing at most of volume-order cannot be worser than that for the Poisson process. Our proofs proceed by bounding the transport or matching cost on large finite windows via the bounds in [13] which are using the expected squared Fourier-transform of the empirical measure of $\mu$. These are strongly related to the scattering intensity, which is known to converge to the Fourier transform of the RPCM, also called the structure factor. These bounds work well for first and second moments but not higher moments. Vanishing of the structure factor at the origin is equivalent to hyperuniformity and quantifying the rate of vanishing on large windows and for small frequencies are the key estimates needed for our proof. Then we transfer these bounds to the infinite sample exploiting tightness and that the finite point processes (i.e., point process on finite windows) converge in law to the infinite sample in the vague topology. This is in contrast to other proofs which use Whitney-type decomposition with error terms estimated via potential theory or PDE methods. Also the other works that derive bounds using the Fourier-analytic method, exploit ’weak independence’ or ’mixing’ of point processes to derive Poisson-like bounds but these do not exploit the vanishing of the structure factor. Our bounds in the finite sample case are explicit and hold for all point processes without assumption of integrable reduced pair correlations. Also in the infinite case, we give explicit bounds for matching when the logarithmic integrability condition is not satisfied by the hyperuniform point processes. Independently, [17] have obtained $p$-Wasserstein transport rates of HU processes on large windows under other additional assumptions on moments and concentration bounds. [32] also recently obtain that HU processes are $L^{2}$ perturbed lattices, assuming finite Coulomb energy, another form of strong hyperuniformity. We give a more detailed comparison with the literature after Theorem 2.
@@
Plan of the paper: In the next subsection, we introduce some preliminaries on point processes. We state our main results and illustrate with examples in Section 2. We also discuss extension of our results to random measures in Section 2.3. The proofs of our main results are in Section 3. In that section, we also state the Fourier-analytic bounds for probability measures due to Bobkov and Ledoux as well as a general proposition that helps us to transfer linear transport cost between large finite samples to finite typical cost for the transport between infinite samples.
@@ 1.1 Preliminaries
[…]
@@
When $\mu,\nu$ are simple point processes, we abusively also denote by $\mu,\nu$ their supports, and we call matching between $\mu,\nu$ a random one-to-one map $T:\nu\to\mu$. If $\nu=\mathcal{L}$, an allocation scheme is a mapping $T:\mathbb{R}^{d}\to\mu$ that is also a transport map, meaning that a.s., $\mathcal{L}(T^{-1}(A))=\nu(A)$ for any $A\subset\mathbb{R}^{d}$. Whether we talk of matching or allocation scheme, we also require that $T$ is invariant (under translations), i.e. the mass transported from $A$ to $B$ measurable subsets of $\mathbb{R}^{d}$ satisfies for a shift $x\in\mathbb{R}^{d}$
@@
$F(r)$ $:=(2\pi)^{-d}\,\mathbf{E}\mathbf{1}(\|T(x)-x\|\leqslant r)\,,$
where $\Lambda_{1}=[-\pi,\pi)^{d}$. To be slightly more general, we shall say that there is an invariant coupling between $\mu$ and $\nu$ whose typical transport distance $X$ has distribution function $F$ if there exists on some probability space a random measure $M$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that $(M(\cdot,\mathbb{R}^{d}),M(\mathbb{R}^{d},\cdot))$ has the same law as $(\mu,\nu)$, and
$F(r)$ $=(2\pi)^{-d}\,\mathbf{E}M(\{(x,y):x\in\Lambda_{1},\|x-y\|\leqslant r\}).$ (2)
This coupling is said a matching if it can a.s. be represented by a one-to-one mapping $T$, i.e. $M=\sum_{x\in\mu}\delta_{(x,T(x))}$. There is also a formulation of typical matching or transport distance in terms of Palm distribution of $\mu$ or $\nu$ but we shall avoid it is not used; see [30, Section 2]. Also when $\mu$ is simple it is easy to see that $\mathbf{E}(w(X))<\infty$ for some function $w:\mathbb{R}_{+}\to\mathbb{R}_{+}$, iff
$\mathbf{E}\sum_{x\in\mu\cap\Lambda_{1}}w(\|x-T(x)\|)<\infty.$
The LHS above can be defined as the cost of transport between $\mu,\nu$ with respect to $w$; see [22, (2.4)]. In the case of $\nu=\mathcal{L}$ or shifted lattice, the above definition yields $X$ with distribution $F$ as mentioned (1) at the beginning of the paragraph for translation-invariant $T$.
@@
[…]
More formally, we assume implicitly that $\mu(A)$ has finite second moment for any bounded set $A$ whenever $\beta$ is mentioned and in which case $\beta$ is a signed measure characterised by the relation
$\mathbf{E}\sum_{x\neq y\in\mu}\varphi(x)\psi(y)$ $=\mathbf{E}\int\int\mathbf{1}(x\neq y)\varphi(x)\psi(y)\mu(\mathrm{d}x)\mu(\mathrm{d}y)$
@@
for non-negative compactly supported continuous measurable functions $\varphi,\psi$. For instance, if $\mu$ is a homogeneous Poisson process, or if $\mu=\mathcal{L}$, $\beta=0.$ Denote by $|\beta|$ the total-variation of $\beta$, and say that $\beta$ is integrable if $|\beta|(\mathbb{R}^{d})<\infty.$ We remark that existence of second moments guarantees that $|\beta|$ is locally integrable, see [38, Section 2].
@@
where $\Lambda_{n}=(-\pi n^{\frac{1}{d}},\pi n^{\frac{1}{d}}]^{d}$. Assuming integrability of RPCM $\beta$, it is known that the choice of window shapes is not relevant i.e., instead of $\Lambda_{n}$ one can consider $n^{1/d}W$ for any convex set $W$ containing the origin in its interior; [38, 57].
@@
when it is well defined. This is the case if for instance the RPCM $\beta$ is integrable. In this situation, HU can be characterized also as $S(0)=0$. Some of these implications hold even without assuming integrability of $\beta$; see [10, Theorem 3.6].

### v4

@@ 1 Introduction
The celebrated AKT (Ajtai, Komlós and Tusnády) theorem [2] states that, if one draws $2n$ i.i.d. uniform points $X_{1},\dots,X_{n},Y_{1},\dots,Y_{n}$ in a two-dimensional cube with volume $n$, it is possible to pair each $X_{i}$ with an exclusive $Y_{\sigma(i)}$ for some permutation $\sigma$ such that the typical length $|X_{1}-Y_{\sigma(1)}|$ is of the order $\sqrt{\ln(n)}$ with high probability, and it is the best possible rate. The role of the dimension $2$ is very important, as for the same statement in dimension $1$, the typical distance is of order $\sqrt{n}$, and in dimension $d\geqslant 3$ it is of order $1.$ This result has known many refinements and generalisations such as replacing the total distance by weighted versions $\sum|X_{i}-Y_{\sigma(i)}|^{p},p>0$, drawing points with a non-uniform law on the cube, giving large deviations [26], computing exact asymptotic constants [5, 34] etc. Using the framework of optimal transport, these asymptotics are equivalent to the cost of optimal transport between the measure $\nu=\sum_{i=1}^{n}\delta_{Y_{i}}$ and the measure $\mu=\sum_{i=1}^{n}\delta_{X_{i}}$ on the cube. We refer the reader to [55, 59, 46] for surveys of such results and for a sampler of other recent developments, see [7, 14, 29].
@@
[…]
@@
The starting point of this article is to replace $\mu$ and/or $\nu$ by an arbitrary stationary random point processes and consider matchings between $\mu$ and $\nu$ and equivalently transport maps or plans between the Lebesgue measure and $\mu$. In specific planar cases such as zeros of Gaussian entire functions (GEF) or Ginibre point process, existence of good matchings have been shown. In particular, [56] shows exponential moments for the matching cost between zeros of GEF and lattice and [38, 52] prove respectively finite first and second moments for the transport of Ginibre ensemble to the Lebesgue measure. Both these point processes are known to be hyperuniform (also known as super-homogeneous) i.e., the asymptotic number variance (i.e. variance of the number of points on a large window) grows more slowly than volume order, unlike the Poisson point process [60, 27]. Hyperuniform (HU) processes appear in various contexts such as statistical physics, biology, material sciences, random matrices, dynamical systems, numerical integration [61, 21]. A concise description coined by physicists about HU processes is “global order and local disorder”, meaning that even though the particles are locally disordered, exhibit isotropy and asymptotic independence, they should behave at large scales as lattices, and should in some sense be arranged in a nice orderly manner. A natural question is therefore whether such processes are well distributed in space, and we believe that this should be reflected by their transport properties, in the sense that a nicely spread point sample should be easy to transport to the Lebesgue measure, or equivalently to a regular lattice. Notable examples of hyperuniform point processes are constructed as a good matching to a lattice; see [51, 41, 47].
@@
We see from (1) that this property (existence of a matching with finite second moment for typical cost) is true in dimension $d\geqslant 3$ for Poisson processes, and we will prove that it extends to other processes, whether HU or not, as long as they are not hyperfluctuating. In dimension $2$, we answer the question on existence of ’good’ matchings for HU processes in the affirmative. In other words, we show that for a HU process $\mu$ in dimension $d=2$, identifying $\mu$ with its support, under mild logarithmic integrability conditions on the reduced pair correlation function, there is a $L^{2}$ process $\{T(k);k\in\mathbb{Z}^{2}\}$ invariant under $\mathbb{Z}^{2}$ translations such that $\mu=\sum_{k\in\mathbb{Z}^{2}}\delta_{k+T(k)}$, meaning it is a $L^{2}$ perturbed lattice. Many known HU processes satisfy this condition, including the Ginibre point process and zeros of Gaussian entire function. An example of a HU process with infinite $L^{1}$-matching has been provided in [35, Theorem 2(2)] indicating the necessity of some additional integrability condition but whether logarithmic integrability is required or not is moot. Our interest in $L^{2}$ perturbed lattices is also in part due to the following recent results on perturbed lattices due to [23]. Namely it was shown that $L^{2}$ perturbed lattices are HU but for any $\delta>0$, there exist $L^{2-\delta}$ perturbed lattices which are not hyperuniform.
@@
En route to our result for HU processes, we generalize the above bounds (1) for Poisson processes to stationary random point processes with integrable reduced pair correlation function (RPCM). This is to say that in magnitude, the matching bounds for point processes with number variance growing at most of volume-order cannot be worser than that for the Poisson process. Our proofs proceed by bounding the transport or matching cost on large finite windows via the bounds in [14] involving the expected squared Fourier-transform of the empirical measure of $\mu$. These are strongly related to the scattering intensity, which is known to converge to the Fourier transform of the RPCM, also called the structure factor. These bounds work well for first and second moments but not higher moments. Vanishing of the structure factor at the origin is equivalent to hyperuniformity and quantifying the rate of vanishing on large windows and for small frequencies are the key estimates needed for our proof. Then we transfer these bounds to the infinite sample exploiting tightness and that the finite point processes (i.e., point process on finite windows) converge in law to the infinite sample in the vague topology. This is in contrast to other proofs which use Whitney-type decomposition with error terms estimated via potential theory or PDE methods. Also the other works that derive bounds using the Fourier-analytic method, exploit ’weak independence’ or ’mixing’ of point processes to derive Poisson-like bounds but these do not exploit the vanishing of the structure factor. Our bounds in the finite sample case are explicit and hold for all point processes without assumption of integrable reduced pair correlations. Also in the infinite case, we give explicit bounds for matching when the logarithmic integrability condition is not satisfied by the hyperuniform point processes. Independently, [18] have obtained $p$-Wasserstein transport rates of HU processes on large windows under other additional assumptions on moments and concentration bounds. [35] also recently obtain that HU processes are $L^{2}$ perturbed lattices, assuming finite Coulomb energy, another form of strong hyperuniformity. We give a more detailed comparison with the literature after Theorem 2.
@@
Plan of the paper: In the next subsection, we introduce some preliminaries on point processes. We state our main results and illustrate with examples in Section 2. We give more detailed proof outlines as well as comparisons with the existing literature. We also discuss the extension of our results to random measures in Section 2.3. The proofs of our main results are in Section 3. In that section, we also state the Fourier-analytic bounds for probability measures due to Bobkov and Ledoux as well as a general proposition that helps us to transfer linear transport cost between large finite samples to finite typical cost for the transport between infinite samples.
@@ 1.1 Preliminaries
A random measure is a random element taking values in the space of locally-finite (or Radon) measures on $\mathbb{R}^{d}$ equipped with the evaluation $\sigma$-algebra. A point process is a random measure taking values in the space $\mathscr{N}$ of locally-finite counting measures on $\mathbb{R}^{d}$ equipped again with the evaluation $\sigma$-algebra. It is simple if a.s., $\mu(\{x\})\in\{0,1\}$ for all $x\in\mathbb{R}^{d}$, and stationary or translation invariant if $\mu+x\overset{d}{=}\mu$ where $(\mu+x)(\cdot):=\mu(\cdot-x)$ is the shifted measure. The set notation is sometimes abusively applied to a simple point process by assimilating it to its support. Common examples of stationary point process are the shifted lattice $\sum_{z\in\mathbb{Z}^{d}}\delta_{z+U}$ for $U$ uniform in $0,1]^{d}$, the homogeneous Poisson process, the infinite Ginibre ensemble, and zeros of Gaussian entire functions; see Section 2.2 or [21, Section 3].
@@
When $\mu,\nu$ are simple point processes, we abusively also denote by $\mu,\nu$ their supports, and we call matching between $\mu,\nu$ a random one-to-one map $T:\nu\to\mu$. If $\nu=\mathcal{L}$, an allocation scheme is a mapping $T:\mathbb{R}^{d}\to\mu$ that is also a transport map, meaning that a.s., $\mathcal{L}(T^{-1}(A))=\nu(A)$ for any $A\subset\mathbb{R}^{d}$. Whether we talk of matching or allocation scheme, we also require that $T$ is invariant (under simultaneous translations), i.e. the mass transported from $A$ to $B$ measurable subsets of $\mathbb{R}^{d}$ satisfies for a shift $x\in\mathbb{R}^{d}$
@@
$F(r)$ $:=(2\pi)^{-d}\,\mathbf{E}\,\nu\{x\in\Lambda_{1}:\|T(x)-x\|\leqslant r\}\,$, (2)
where $\Lambda_{1}=(-\pi,\pi]^{d}$. To be slightly more general, we shall say that there is an invariant coupling between $\mu$ and $\nu$ whose typical transport distance $X$ has distribution function $F$ if there exists on some probability space a random measure $M$ on $\mathbb{R}^{d}\times\mathbb{R}^{d}$ such that $(M(\cdot,\mathbb{R}^{d}),M(\mathbb{R}^{d},\cdot))$ has the same law as $(\mu,\nu)$, and
$F(r)$ $=(2\pi)^{-d}\,\mathbf{E}M(\{(x,y):x\in\Lambda_{1},\|x-y\|\leqslant r\}),$ (3)
and the random measure $M$ is invariant under simultanenous translations or equivalently equivariant, i.e. $M(x+\cdot,x+\cdot)\stackrel{{\scriptstyle(d)}}{{=}}M$ for all $x\in\mathbb{R}^{d}$. This coupling is a matching if it can a.s. be represented by a one-to-one mapping $T$, i.e. $M=\sum_{x\in\nu^{\prime}}\delta_{(T(x),x)}$ a.s., where $\nu^{\prime}=M(\mathbb{R}^{d},\cdot).$ Observe that matchings induced by invariant couplings are translation-invariant.
There is also a formulation of typical matching or transport distance in terms of Palm distribution of $\mu$ or $\nu$ but we do not require it here; see [33, Section 2]. Also when $\mu$ is simple it is easy to see that $\mathbf{E}(w(X))<\infty$ for some function $w:\mathbb{R}_{+}\to\mathbb{R}_{+}$ iff
$\mathbf{E}\sum_{x\in\nu^{\prime}\cap\Lambda_{1}}w(\|x-T(x)\|)<\infty.$ (4)
The LHS above can be defined as the cost of transport between $\mu,\nu$ with respect to $w$; see [25, (2.4)].
@@
$\beta(dx)=\mathbf{P}(\mathrm{d}x\in\mu\;|\;0\in\mu)-\mathbf{P}(dx\in\mu).$
More formally, we assume implicitly that $\mu(A)$ has finite second moment for any bounded set $A$ whenever $\beta$ is mentioned, in which case $\beta$ is a signed measure over compact sets characterised by the relation
$\mathbf{E}\sum_{x\neq y\in\mu}\varphi(x)\psi(y)$ $={\mathbf{E}\int\int\mathbf{1}[x\neq y]\varphi(x)\psi(y)\mu(\mathrm{d}x)\mu(\mathrm{d}y)}$
@@
for non-negative compactly supported continuous functions $\varphi,\psi$. For instance, if $\mu$ is a homogeneous Poisson process, or if $\mu=\mathcal{L}$, then $\beta=0.$ Denote by $|\beta|$ the total-variation of $\beta$, and say that $\beta$ is integrable if $|\beta|(\mathbb{R}^{d})<\infty.$ We remark that existence of second moments guarantees that $|\beta|$ is locally integrable and furthermore stationarity of $\mu$ implies that $\beta$ is an even function i.e., $\beta(\mathrm{d}x)=\beta(-\mathrm{d}x)$; for example, see [22, Section 8] for more on second order properties of random measures.
@@
where $\Lambda_{n}=(-\pi n^{\frac{1}{d}},\pi n^{\frac{1}{d}}]^{d}$. Assuming integrability of the RPCM $\beta$, it is known that the choice of window shapes is not relevant i.e., instead of $\Lambda_{n}$ one can consider $n^{1/d}W$ for any convex set $W$ containing the origin in its interior; for example, see [21, Proposition 2.2].
@@
when it is well defined. This is the case if for instance the RPCM $\beta$ is integrable. In this situation, HU can be characterized also as $S(0)=0$. Some of these implications hold even without assuming integrability of $\beta$; see [11, Theorem 3.6].
Let us finally give a toy example of an HU process with an integrable RPCM which is also a perturbed lattice; to our knowledge it was first emphasized in [40]. It is called cloaked peturbed lattice, denoted by $\nu_{\textrm{Cl}}$, because the periodic component is hidden at the second order, or cloaked, by the vanishing of the Fourier transform of the $U_{k}$ on the grid. It helps make a bridge between disordered systems such as Poisson or determinantal, and perturbed lattices.
Example 1.
Consider the shifted lattice perturbed by i.i.d variables $U_{k},k\in\mathbb{Z}^{d}$ uniform in $0,1]^{d}$
$\nu_{\textrm{Cl}}:=\{k+U+U_{k};k\in\mathbb{Z}\},$
where $U$ is also a uniform random variable in $0,1]^{d}$ and independent of the $U_{k}$’s. For $\varphi,\psi$ as above,
$\mathbf{E}\sum_{x\neq y\in\nu_{\textrm{Cl}}}\varphi(x)\psi(y)=$ $\sum_{k\neq m\in\mathbb{Z}^{d}}\int_{([0,1]^{d})^{3}}\varphi(k+u+v)\psi(m+u+w)\,\mathrm{d}u\,\mathrm{d}v\,\mathrm{d}w$
$=$ $\int_{[0,1]^{d}}\sum_{k,m\in\mathbb{Z}^{d}}\left(\int_{[0,1]^{d}}\varphi(k+u+v)\,\mathrm{d}v\int_{[0,1]^{d}}\psi(m+u+w)\mathrm{d}w\right)\mathrm{d}u$
$\hskip 56.9055pt-\int_{[0,1]^{d}}\,\sum_{k\in\mathbb{Z}^{d}}\left(\int_{([0,1]^{d})^{2}}\varphi(k+u+v)\,\psi(k+u+w)\,\mathrm{d}v\,\mathrm{d}w\right)du$
${=\int_{\mathbb{R}^{d}}\,\varphi\int_{\mathbb{R}^{d}}\psi-\int_{[0,1]^{d}}\sum_{k\in\mathbb{Z}^{d}}\left(\int_{([0,1]^{d})^{2}}\,\varphi(k+u+v)\,\psi(k+u+w)\mathrm{d}v\,\mathrm{d}w\right)du}.$
Swapping the order of summation and integrals, the second term can be re-written as
$\int_{[0,1]^{d}}\sum_{k\in\mathbb{Z}^{d}}\left(\int_{([0,1]^{d})^{2}}\,\varphi(k+u+v)\,\psi(k+u+w)\mathrm{d}v\,\mathrm{d}w\right)du$
$=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(z+v)\,\psi(z+v+w)\,\mathbf{1}[v\in[0,1]^{d},v+w\in[0,1]^{d}]\mathrm{d}z\,\mathrm{d}v\,\mathrm{d}w$
$=\int_{\mathbb{R}^{d}\times\mathbb{R}^{d}}\varphi(z)\,\psi(z+w)\,\mathcal{L}([0,1]^{d}\cap([0,1]^{d}-w))\,\mathrm{d}z\,\mathrm{d}w,$
and so from the characterisation (5), we see that the RPCM of the i.i.d. perturbed lattice $\nu_{\textrm{Cl}}$ is $\beta(\mathrm{d}x)=\mathcal{L}([0,1]^{d}\cap([0,1]^{d}-x))\,\mathrm{d}x$. Thus the RPCM is supported by $B(0,2\sqrt{d})$ and clearly integrable. Furthermore, the obvious matching $T(k)=k+U+U_{k},k\in\mathbb{Z}^{d}$ from $\mathbb{Z}^{d}$ to $\mu$ has typical matching distance $X\leq 2\sqrt{d}$ a.s.. Also, we know that $\nu_{\textrm{Cl}}$ is hyperuniform; for instance, see [10].